OpenStudy (anonymous):

Which expression is not equivalent to the other expressions? seven plus zero point five times four raised to the second power two point five times the quantity sixteen minus two raised to the second power, end quantity the quantity three raised to the third power plus three end quantity divided by two the quantity twenty plus five raised to the second power, end quantity, divided by three

OpenStudy (anonymous):

I may have made a mistake, but these are the answers I have. 7+.5(4^2)=15 2.5(16-2^2)=30 (3^3+3)/2=6 (20+5^2)/3=15

OpenStudy (anonymous):

I found my mistake, (3^3+3)/2=15 sorry.